OpenStudy (anonymous):

A customer has an offer for 5 portrait sheets for 9.99 with additional portrait sheets priced at 19.99 each. If the customer desires to purchase 10 portrait sheets, how much would she save if she instead purchases a 10 portrait sheet package for $99.99?

OpenStudy (anonymous):

5 * 19.99 = 99.95 + 9.99 = 109.94 109.94 - 99.99 = 9.95 The customer would save $9.95.

OpenStudy (anonymous):

The question states "5 portrait sheets for $9.99 with additional portrait sheets priced at $19.99 each." The first five of those portrait sheets all come together for $9.99. If not, the question is worded improperly.